OpenStudy (anonymous):

Identify intervals on which the function is increasing, decreasing, or constant. g(x) = 1 - (x - 7)^2

OpenStudy (anonymous):

are you in a calculus class?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok, then you probably want to find the second derivative first, because there is a second derivative test which will tell you if the function is increasing or decreasing.

OpenStudy (anonymous):

okay.. then what?

OpenStudy (anonymous):

I figured it out

OpenStudy (anonymous):

g(x) = 1 - (x - 7)^2 d/dx(g(x)) = -2(x-7) d/dx(-2(x-7)) = -2 http://en.wikipedia.org/wiki/Second_derivative_test if the second derivative is less than 0 then the function is concave down. since this is a parabola, it is always concave down as seen in this graph. http://www.wolframalpha.com/input/?i=g%28x%29+%3D+1+-+%28x+-+7%29%5E2 however, you are asking when it is increasing and where it is decreasing for that we need the first derivative. set it equal to zero to find the x value where the function is a maximum -2(x-7) = 0 -2x + 14 = 0 -2x = -14 x = 7 the function is increasing from (-infinity,7) and decreasing from (7 to infinity). a good video to learn more about this: http://www.youtube.com/watch?v=g8Nv7Gap2Z8